OpenStudy (anonymous):

f(x) = 1-f(x-1) For positive integer x, if f(2)=12, then compute f(2006).

OpenStudy (anonymous):

Oh god lol

myininaya (myininaya):

oops that x=2 one sec

myininaya (myininaya):

\[f(2)=1-f(2-1)=1-f(1)=12=> f(1)=-11\]

myininaya (myininaya):

\[f(3)=1-f(2)=1-12=-11\] \[f(4)=1-f(3)=1-(-11)=12\] \[f(5)=1-f(4)=1-12=-11\]

OpenStudy (anonymous):

Thank you! But what about f(2006)? I'm confused on how to incorporate that..

myininaya (myininaya):

i think there may be a pattern

OpenStudy (anonymous):

I see..

myininaya (myininaya):

\[f(6)=1-f(5)=1-(-11)=12\] \[f(7)=1-f(6)=1-12=-11\]

myininaya (myininaya):

see its either -11 or 12

OpenStudy (anonymous):

all odds are -11 evens are 12

myininaya (myininaya):

right

OpenStudy (anonymous):

Wow, thanks so much you guys! First time using this site and I love it already. (: